### v1

# Hypertree shrinking avoiding low degree vertices

Abstract
The shrinking operation converts a hypergraph into a graph by choosing, from each hyperedge, two endvertices of a corresponding graph edge. A hypertree is a hypergraph which can be shrunk to a tree on the same vertex set. Klimošová and Thomassé [J. Combin. Theory Ser. B 156 (2022), 250–293] proved (as a tool to obtain their main result on edge-decompositions of graphs into paths of equal length) that any rank $3$ hypertree $T$ can be shrunk to a tree where the degree of each vertex is at least $1/100$ times its degree in $T$. We prove a stronger and a more general bound, replacing the constant $1/100$ with $1/2k$ when the rank is $k$. In place of entropy compression (used by Klimošová and Thomassé), we use a hypergraph orientation lemma combined with a characterisation of edge-coloured graphs admitting rainbow spanning trees.
11footnotetext: Department of Mathematics and European Centre of Excellence NTIS (New Technologies for the Information Society), University of West Bohemia, Pilsen, Czech Republic.22footnotetext: E-mail: khylas@kma.zcu.cz.33footnotetext: E-mail: kaisert@kma.zcu.cz.
1 Introduction
It is well known that each graph has an associated matroid. Lorea [Lor] extended this association from graphs to hypergraphs; the bases of the matroid associated with a hypergraph $H^{\prime}$ are the hypertrees in $H^{\prime}$, i.e., subhypergraphs of $H^{\prime}$ satisfying the following definition. (We refer to Section 2 for any other necessary definitions.)
A hypergraph $H$ is a hypertree if it is possible to choose two vertices from each hyperedge in such a way that the chosen pairs, viewed as edges of a graph on vertex set $V(H)$, form a tree. An equivalent definition may be given in terms of a Hall-type condition (cf. [FKKr, Theorem 2.2]). The equivalence was proved by Lovász [Lov].
The above operation that produces a graph from a hypergraph by choosing a pair of vertices in each hyperedge will be called shrinking. Thus, a hypertree is a hypergraph which can be shrunk to a tree.
Klimošová and Thomassé [klimosova] derived a result on hypertree shrinking that preserves vertex degree up to a constant factor, and used it as one of the tools needed to obtain their main result about decompositions of 3-edge-connected graphs into paths of equal length. Their lemma on shrinking [klimosova, Lemma 22] is as follows:
Lemma 1.
Let $H$ be a hypertree with hyperedges of size at most three. It is possible to shrink $H$ to a tree $T$ such that for every vertex $v$ of $H$,
$d_{T}(v)\geq\frac{d_{H}(v)}{100}.$
Lemma 1 is proved using entropy compression, the method devised to prove an algorithmic version of the Lovász Local Lemma [MT].
In this paper, we use a different method to strengthen Lemma 1 in two ways: first, our version ensures a stronger degree bound, and second, it applies to hypergraphs with arbitrary size of the hyperedges. In addition, the proof is conceptually simpler. We prove:
Theorem 2.
Let $H$ be a hypertree with hyperedges of size at most $k$. It is possible to shrink $H$ to a tree $T$ such that for every vertex $v$ of $H$,
$d_{T}(v)\geq\frac{d_{H}(v)}{2k}.$
2 Preliminaries
We review some of the basic definitions on hypergraphs. A hypergraph is a pair $H=(V,E),$ where $V$ is a finite set and $E$ is a set of subsets of $V$. The elements of $V$ are the vertices of $H$ and the elements of $E$ are the hyperedges of $H.$
The rank of a hypergraph $H$ is the maximum size of a hyperedge of $H$. If all hyperedges of $H$ have size exactly $k$, then $H$ is $k$-uniform. The number of hyperedges of $H$ containing a vertex $v\in V$ is the degree of $v$ and is denoted by $d_{H}(v).$
A hypergraph is simple if it does not contain loops (i.e., hyperedges of size 1) nor repeated hyperedges (distinct hyperedges with identical vertex sets). All the hypergraphs discussed in this paper will be simple (and we will henceforth drop this adjective). A hypergraph is connected if for every nonempty proper subset $X$ of its vertex set, there is a hyperedge intersecting both $X$ and its complement.
It will be useful in our argument to consider directed hypergraphs, consisting of a set of hyperarcs on a finite vertex set. A hyperarc is a hyperedge $e$ together with a designated head. The other vertices of $e$ are called the tails of $e$.
The indegree of vertex $v$ in a directed hypergraph $\vec{H}$, denoted by $d_{\vec{H}}^{IN}(v)$, is the number of hyperarcs whose head is $v$. Similarly, the outdegree of $v$ (denoted by $d_{\vec{H}}^{OUT}(v)$) is the number of hyperarcs in which $v$ is a tail. Observe that a hyperarc of size $k$ contributes to the indegree of exactly one vertex and to the outdegree of $k-1$ vertices.
3 Tools
In the proof of Theorem 2, we will need two main tools. The first one is an orientation lemma for hypergraphs with lower bounds on the indegrees. The second one is a characterisation of edge-coloured graphs admitting rainbow spanning trees, described in the last few paragraphs of this section. Throughout this section we define $\sum_{v\in X}f(v)$ as $f(X).$
We begin with the topic of hypergraph orientation. A result of Frank, Királyi and Királyi [FKK, Lemma 3.3] gives a necessary and sufficient condition for the existence of an orientation of a hypergraph with prescribed indegrees. We use it in an equivalent formulation from [Ramsey].
Lemma 3 ([FKK]).
Let $H=(V,E)$ be a hypergraph and let $f:V\rightarrow\mathbb{Z}^{+}$ be a mapping of the vertex set $V$ of $H$ into the set of non-negative integers. Then there is an orientation $\vec{H}$ of $H$ such that $d^{IN}_{\vec{H}}(v)=f(v)$ for every $v\in V$ if and only if
  1. (i)
    $f(V)=e(H),$
  2. (ii)
    $f(F)\geq e(F)$ for every $F\subseteq V,$
where $e(F)$ denotes the number of hyperedges in the subset $F\subseteq V$ and $e(H)$ is the number of all hyperedges in $H$.
What we actually need is a sufficient condition for the existence of an orientation where the indegree of each vertex satisfies a given lower bound. This is similar to [Ramsey, Lemma 3], except that upper bounds are prescribed in the latter lemma. [^1] It may seem that in [Ramsey], Lemma 3 is formulated for $d^{OUT}_{\vec{H}}(v)$ rather than $d^{IN}_{\vec{H}}(v)$; this is not the case since the definitions used in [Ramsey] exchange the notions of outdegree and indegree. For hypergraphs (as opposed to graphs), a version with lower bounds does not seem to be a trivial consequence of one with upper bounds, but it turns out to be possible to follow the proof from [Ramsey] with suitable modifications. Below, we give the proof in full detail to make this paper self-contained.
Lemma 4.
Let $H=(V,E)$ be a hypergraph and let $f:V\rightarrow\mathbb{Z}^{+}$ be a mapping of the vertex set $V$ of $H$ into the set of non-negative integers. Assume that for every $F\subseteq V$,
$f(F)\leq e^{*}(F),$ (1)
where $e^{*}(F)$ denotes the number of hyperedges incident with $F$. Then there is an orientation $\vec{H}$ of $H$ such that
$d^{IN}_{\vec{H}}(v)\geq f(v)$ (2)
for every $v\in V.$
Proof.
Let $g:V\rightarrow\mathbb{Z}^{+}$ be a mapping such that
  1. (i)
    $g(F)\leq e^{*}(F)$ for every $F\subseteq V,$
  2. (ii)
    $g(v)\geq f(v)$ for every $v\in V,$
  3. (iii)
    $g(V)$ is the maximum among all functions that satisfy (i) and (ii).
We will verify properties (i) and (ii) of Lemma 3 for the function $g$. In other words, we will show that
$g(V)=e^{*}(H)=e(H),$ (3)
where $e(H)$ denotes the number of hyperedges of the hypergraph $H$, and that $g(F)\geq e(F)$ for every $F\subseteq V.$ Then, by Lemma 3, there is an orientation $\vec{H}$ of $H$ for which $d^{IN}_{\vec{H}}(v)=g(v)\geq f(v)$ for every $v\in V.$
Let $X\subseteq V$ be a set with maximum cardinality for which $g(X)=e^{*}(X).$ Firstly, we want to show that $V\backslash X$ does not contain any vertex $w$ with the function $g(w)$ equal to the degree of $w$ in the hypergraph $H.$
It is possible that $X$ is an empty set. If $X=V$ then we are done, so assume that $|X|<|V|.$ If $w$ is a vertex in $V\backslash X$ for which $g(w)=d_{H}(v),$ then by the maximality of $|X|$ we obtain
$\begin{split}g(X\cup\{w\})&<e^{*}(X\cup\{w\})\leq e^{*}(X)+e^{*}(\{w\})=\\ &=g(X)+d_{H}(w)=g(X)+g(w)=g(X\cup\{w\}),\end{split}$
which is a contradiction. Therefore $V\backslash X$ contains no such vertex $w.$
Furthermore, we will prove that there exists a function $h(v)$ which satisfies properties (i) and (ii), but for every vertex $v\in V$ it holds that $h(v)>g(v)$ which contradicts the maximality of $g(V).$
Let $z$ be a vertex in $V\backslash X$ and let $h:v\rightarrow\mathbb{Z}^{+}$ be a mapping such that $h(v)=g(v)$ for all $v\in V\backslash\{z\}$ and $h(z)=g(z)+1.$ Suppose that there exists a subset $F\subseteq V$ such that $h(F)>e^{*}(F).$ Then $z\in F$ and
$e^{*}(F)\geq g(F)=h(F)-1>e^{*}(F)-1,$
which implies that $g(F)=e^{*}(F).$ But by the maximality of $|X|$ and because $F$ is not contained in $X,$ as $z\in F$ and $z\in V\backslash X,$ we get that
[…]
which is again a contradiction. [^2] We would like to mention that in the corresponding part of the proof in [Ramsey], the inequality analogous to (4) is slightly incomplete, but this does not affect the validity of the subsequent arguments.
Therefore, for all subsets $F\subseteq V$ it holds that $h(F)\leq e^{*}(F).$ Also for every vertex $v\in V$ it is true that $h(v)\geq g(v)\geq f(v),$ so the function $h$ satisfies properties (i) and (ii). However, $h(V)=g(V)+1,$ i.e. $h(V)>g(V)$ which is a contradiction with the maximality of $g(V).$ Hence, property (i) from Lemma 3 holds.
Let $F^{\prime}=V\backslash F$ for every $F\subseteq V.$ We know that $g(V)=e(H)$ and it is also true that
$g(V)-g(F)=e(H)-g(F)$
Therefore,
$e(H)-g(F)=g(F^{\prime})\leq e^{*}(F^{\prime})=e(H)-g(F).$
From this we get that $g(F)\geq e(F)$ for every $F\subseteq V,$ i.e., property (ii) from Lemma 3 holds as well. ∎
The second main tool which we will use to prove Theorem 2 is a necessary and sufficient condition for an edge-coloured graph to contain a rainbow spanning tree.
Let $G$ be a graph with a (not necessarily proper) edge colouring. A subgraph of $G$ is rainbow if it does not contain two edges with the same color. The following necessary and sufficient condition for the existence of a rainbow spanning tree has been derived in [heterochromatic_ST] and, independently, in [Schrijver, Section 41.1a].
Theorem 5 ([heterochromatic_ST]).
Let $G$ be a (possibly improperly) edge-coloured graph of order $n$. There exists a rainbow spanning tree of $G$ if and only if
for any set of r colours (0rn2), the removal of all edges colouredwith these r colours from G results in a graph with at most $\begin{split}&\textrm{for any set of $r$ colours $(0\leq r\leq n-2)$, the removal of all edges coloured}\\ &\textrm{with these $r$ colours from $G$ results in a graph with at most $r+1$}\\ &\textrm{components.}\end{split}$ (5)
Observe that the case $r=0$ of condition (5) corresponds to $G$ being connected.
4 Shrinking hypertrees
In this section, we prove Theorem 2. We begin with an application of Lemma 4 concerning a specific lower bound for the indegrees in an orientation of a hypergraph.
Lemma 6.
Let $H=(V,E)$ be a hypergraph of rank at most $k$. There exists an orientation $\vec{H}$ of $H$ in which
$d^{IN}_{\vec{H}}(v)\geq\left\lfloor\frac{d_{H}(v)}{k}\right\rfloor$
for every vertex $v\in V.$
Proof.
We apply Lemma 4 with $f(v)=\lfloor d_{H}(v)/k\rfloor$. For a set $F\subseteq V$ of vertices of $H$, we verify condition (1). Clearly,
$\sum_{v\in F}f(v)\leq\sum_{v\in F}\frac{d_{H}(v)}{k}.$ (6)
Each hyperedge $e$ of $H$ contributes at most $1$ to the degree of a specific vertex $v$, and since $|e|\leq k$, its total contribution to the right hand side of (6) is at most $k\cdot\frac{1}{k}=1$. It follows that the right hand side is at most $e^{*}(F)$, i.e., (1) holds. ∎
Let $H$ be a hypertree of rank $k$, and let $\vec{H}$ be the orientation from Lemma 6. We construct an edge-coloured graph $G(\vec{H})$ on the vertex set of $\vec{H}$ by the following rule: for each hyperarc $\vec{e}$ of $\vec{H}$, add to $G$ a star whose center is the head of $e$ and whose leaves are the tails of $e$. Furthermore, all the edges of this star have the same colour, which differs from the colours used for the other stars.
We claim that $G(\vec{H})$ has a rainbow spanning tree. To prove this, we need to verify condition (5) of Theorem 5. Consider first a simple variant of the above construction: let the graph $G_{H}$ be obtained by adding, for each hyperedge $e$ of $H$, a complete graph on the vertex set of $e$. Observe that since $H$ is a hypertree, $G_{H}$ has a rainbow spanning tree. Consequently, condition (5) holds for $G_{H}$.
Now in the construction of $G(\vec{H})$, we used stars in place of the complete subgraphs. It is, however, easy to see that this replacement has no effect on the validity of condition (5): indeed, if instead of each complete subgraph we take any connected subgraph on the same vertex set, then their union has the same number of components in each case, and so the validity of condition (5) remains without change.
Thus, $G(\vec{H})$ has a rainbow spanning tree $T^{\prime}$ as well. Note that for each hyperedge $e$ of $H$, $T^{\prime}$ has one edge $e^{\prime}$ corresponding to $e$, namely the edge selected from the star corresponding to an orientation $\vec{e}$ of $e$. Moreover, the head of $\vec{e}$ is one of the endvertices of $e^{\prime}$.
By the above, if $v$ is a vertex of $H$, then the degree of $v$ in $T^{\prime}$ is at least the indegree of $v$ in $\vec{H}$ — that is, at least $\lfloor d_{H}(v)/k\rfloor$. If $d_{H}(v)\geq k$, then we have
$d_{T^{\prime}}(v)\geq\frac{d_{H}(v)}{2k}$ (7)
since, clearly, $\left\lfloor x\right\rfloor\geq x/2$ for all real $x\geq 1$. On the other hand, if $d_{H}(v)<k$, then (7) holds as well, because the right hand side is smaller than $1/2$ while the degree of $v$ in the tree $T^{\prime}$ is at least $1$. Thus, (7) holds for all vertices, and we have found the required shrinking of $H$. This concludes the proof of Theorem 2.
[…]